A committee of 4 women and 3 men need to select a chair person and a secretary. If every person is equally likely to be chosen, what is the probability that :
a) both positions are filled by women?
b) both positions are filled by men?
c) the chairperson is female and the secretary is male?

Answer by stanbon(75887)

A committee of 4 women and 3 men need to select a chair person and a secretary. If every person is equally likely to be chosen, what is the probability that :
a) both positions are filled by women?
Ans: 4C2/7C2 = 6/21 = 2/7
------------------
b) both positions are filled by men?
Ans: 3C2/7C2 = 3/21 = 1/7
-------------------------------
c) the chairperson is female and the secretary is male?
# of ways to succeed: 4*3 = 12
# of possible outcomes: 7*6
Prob = 12/42 = 6/21 = 2/7
------------------------------
Cheers,
Stan H.